OpenStudy (abbycross167):

Can someone please help me to answer this math question? Please? I will fan and give a medal...

OpenStudy (will.h):

sure

OpenStudy (abbycross167):

A tank can be filled by one pipe in 9 minutes and drained by another pipe in 12 minutes. Determine how long it will take the fill the tank

OpenStudy (will.h):

do you need to form a linear function? or do you only need the time?

OpenStudy (abbycross167):

I'm sure just the time but I have to show the steps

OpenStudy (will.h):

1/9 - 1/12 = 1/x therefore x= 1/36 however i think you should change the places therefore the answer is 36 mins hope that helps... somehow

OpenStudy (mathmate):

The way that works is that if it fills in 9 hours, that means in one hour it fills 1/9 of the tank. Same goes for the drain. It drains in 12 hours means it drains 1/12 of the tank in one hour. So in one hour it will fill a net of 1/9-1/12=1/36 of the tank.Therefore it takes 36 hours to fill the tank with the drain unplugged.

OpenStudy (mathmate):

Correction: everything is in minutes! sorry.

OpenStudy (will.h):

There's another way... Let x=time it takes for tank to overflow when both pipes are open and the tank is empty When both pipes are open the tank fills at the rate of 1/x tank per min Inlet pipe fills at the rate of 1/9 tank per min drain empties at the rate of 1/12 tank per min So, our equation to solve is: 1/9 - 1/12=1/x multiply each term by 60x

OpenStudy (abbycross167):

Thank you both so much!!! y'all are very helpful!!